Please on supply function and inverse supply function. If you are given the equation: Qx =-30+2Px_4Pz , suppose Pz is $60 , what is the supply function and the inverse supply function.


1
Expert's answer
2021-03-08T09:22:16-0500

(a) If "P_z=\$60", then the supply function for good X takes form:


"Q_x^s=-30+2P_x-4\cdot60=-270+2P_x."

(b) Let's find the inverse supply function:


"Q_x^s+270=2P_x,""P_x=0.5Q_x^s+135."